You are to start a Dopamine drip at 2.5 mcg/kg/min for renal perfusion. The patient weighs 150 lbs. The stock from the pharmacy reads Dopamine 400 mg/500 mL D5W.

What is the flow rate in mL/hr?
Answer #1

Orderd Dopamine drip at 2.5 mcg/kg/min

Patient weighs= 150 lbs= 150/2.2kg= 68.18 kg

As per weight of the patient required dose= 2.5 X 68.18=170.45 mcg/min

Available Dopamine= 400mg/500ml

=0.8 mg/ml= 800mcg/ml

800mcg in 1 ml

1 mcg in= 1/800 ml

170.45 mcg in = 170.45/800 ml=0.213 ml

Hence 170.45 mcg/min= 0.213 ml/min

= 12.78 ml/hr is the required flow rate in ml/hr.